a bank teller has some 5 dollar bills and some twenty dollar bills. The teller has 10 more of the 20's. The total value of the moey is 775. how many 5's are there?
Answer by mukhopadhyay(490)

Assume that the teller has x number of $5 bills.
=> The teller has x+10 $20 bills
5x + (x+10)(20) = 775
=> 5x+20x+200 = 775
=>x=575/25=23
Answer: The teller has 23 5's.